A rectangular field is to be enclosed by a fence then subdivided into three areas by fences

parallel to the shorter side. Find the dimensions of the field with the largest total area that can be

enclosed with 260 m of fencing. What is the area?


1
Expert's answer
2022-04-12T15:55:54-0400

As rectangular field is divided into three areas the fence is consist of 4 short sides and 2 large sides.

Define "x" m is a short side. Then "(260-4x)\/2" m is a large side. Total area:

"S(x)=x(260-4x)\/2"

Let’s find x for which the total area is maximum:

"S\u2019(x)=130-4x=0"

"x=32.5"

As "S\u2019(30)=130-120>0" and "S\u2019(35)=130-140<0" we can conclude "x=32.5" corresponds to maximum value "S(x)".

Sides of the rectangular field:

32.5 m and "(260-4\cdot32.5)/2=65" m.

Largest total area:

"S=32.5\cdot65=2112.5" "m^2"